# Graph Reinforcement Learning in Power Grids: A Survey

### v1

@@ 3.1.1 Transmission grids
Transmission grids and their assets operate at high voltages ranging from 110 kV up to 765 kV. The exact voltage level is country-specific, therefore a clear distinction of voltage levels cannot be made. For example, in Germany the 110 kV network is considered as the high-voltage distribution grid. However, the purpose of transmission grids is across all countries the same. It aims to transport large quantities of electrical power over a large distances. The distances vary from a few hundred kilometers (e.g., connecting off-shore wind in Northern Germany with the West-German industry) to a few thousands (e.g., China’s cross-country interconnections).
Transmission grid operation aims for at minimum N-1 secure operation, i.e., of one asset in the grid fails, the remaining grid is not overloaded. Therefore, transmission grids are typically built in meshed structures with a lot of redundancies installed, e.g, multiple transformers and busbars at substations. The structure of transmission grids results in a highly-complex system where optimization problems have to be solved on a large scale with robust results.
One remedial action to prevent grid congestion, e.g., transformer or line overloading is re-dispatch which refers to the change of generator injections. As generation and consumption in power system have to be balanced at all times, the change of one generator set point has to result in a set point change of another generator. This can result in a step down of renewable generators and a step up of fossil-fueled generators, which is an undesirable action, both in terms of costs and CO2-neutrality. Therefore, other means of flexibility are researched such as topology control of the transmission system. Controlling the switching state in substations, the topology of the grid can be modified and help in decreasing or even completely avoiding the need for re-dispatch. Considering the topology into optimization is a challenge itself, as it results in a Mixed-Integer Non-Linear Problem. Here, AI solutions such as RL can help and research on these challenges are on-going Marot et al., (2020).
@@ 3.1.2 Distribution grids
Distribution grids are operated on lower voltage levels, typically lower than 110 kV (e.g., 12.47 kV and 20 kV) until voltage levels of e.g., 120 V and 230 V. The purpose of distribution grids is to deliver the transmitted electrical power from the transmission grids to its individual customers, such as industrial consumers at medium-voltage level and directly to households.
The geographical area covered by distribution grids is smaller than for transmission grids, e.g., one low-voltage grid serving a village with electricity. As the impact of asset failures is rather low in comparison to transmission grids, distribution grids are often not designed considering the N-1 security criterion. Therefore, radial or open-loop structures are common in distribution grid operation.
This results in a higher volatility to voltage fluctuations due to changing generation or consumption over time. With up-ramping installations of photovoltaic generation, this voltage fluctuations can cause voltages to exceed the limits set by the respective DSO. Traditionally, regulated distribution transformers, shunt capacitors, and voltage regulators are used for voltage control Srivastava et al., (2023). With ongoing digitization efforts, the possibilities for voltage control are increasing. Inverter-based technologies, such as smart PV inverters Howlader et al., (2020), vehicle-to-grid Gonzalez Venegas et al., (2021), and stationary batteries Stecca et al., (2020) are flexibilities that become necessary to be tapped by grid operators to maintain the system in a secure state. Due to the high number of distribution grids in the power system, optimized control strategies need scalable and robust solutions, where RL-based solutions can help.
@@ 3.1.3 Grid models
In research there are multiple instances of synthetic grid models. Their purpose is to have a common benchmark for researchers across power system engineering to test new concepts and to be able to compare results with each other. There are multiple examples in literature for benchmark grid models Babaeinejadsarookolaee et al., (2021); Meinecke et al., 2020a ; Meinecke et al., 2020b ; Strunz et al., (2014). In the above mentioned use cases of transmission and distribution system control, the IEEE test cases are the most commonly used grid benchmarks, available in multiple power system calculation tools like MATPOWER Zimmerman et al., (2011), pandapower Thurner et al., (2018), and DIgSILENT PowerFactory DIgSILENT GmbH, (2024).
However, these benchmark cases are simplified cases of power systems which do not fully capture the complexity and challenges of power system operation. These grid models are mostly simple bus branch models with loads and generators connected to the buses. More sophisticated assets like, inverter-based generator controls, transformer tap changers, shunt elements, etc., are not common for such grid models but are common practice in real-life power system operation. Therefore, methods benchmarked on such synthetic grid models have to be treated cautiously as the application in practice requires further considerations.
@@ 3.2 Graph Neural Networks
Figure 2: Left: Visualization of the general message passing scheme in GNNs - The target node (orange) receive messages $m_{u}i$ from its neighbours and aggregates them. The messages can be constructed from information of both, the target and neighbouring node depending on the message passing scheme. Right: Illustration of a GNN - The graph is input to the GNN layers which compute node embeddings based on the messages from neighbouring nodes. As indicated in orange, this is done for each node in the graph. After all embeddings are computed, an activation function is applied. This is repeated for a given number of layers. In the end, the GNN outputs a graph with new node feature from which prediction can be made.
GNNs are designed to extract information from graph-structured data by applying multiple layers of graph convolutions. They can be interpreted as a generalisation of Convolutional Neural Networkss to non-euclidean structured data. The general idea is to combine information from local regions of the inputs in a learnable way and to grow these local regions from layer to layer. In this way, CNNs learn increasingly abstract features from the input data from layer to layer. CNNs perform extremely well on grid-structured data, such as images. Here, the neighbourhood of a pixel consists of the neighbouring pixels, forming a regular grid. However, many real-world phenomena involve relationships or complex dependencies that cannot be represented as regular grid structures without losing information. For example, in a regular grid every node has the same number of neighbours, but in power grids not every components is connected to the same number of power lines.
Graphs consist of an unordered set of nodes and edges, where the edges define the neighbourhood of a node. Graphs can therefore be used to model complex relationships, such as neighbourhoods of arbitrary size or multiple types of edges or nodes. They can have attributes that describe properties of nodes and edges, such as node features or the strength of an edge. This makes graphs a perfect model for many real-world applications, such as power grids. In addition, feed-forward neural networks working on Euclidean data typically treat nodes as independent samples. This means, they neglect the relationships between nodes, or stack them unsystematically into a vector. GNNs, on the contrary, make use of the information about node connectivity. They can solve all common learning tasks, i.e. classification, regression, and clustering, for entire graphs, and at node- or edge level.
@@ 3.2.1 Message Passing
GNNs update the embeddings of the graph nodes by repeatedly aggregating information of their neighbourhoods in a learnable way. This general scheme is called message passing as each node updates its embedding based on the messages it receives from its neighbours. The representation $\bm{h}_{u}$ of a target node $u$ generated by a general message passing layer is computed as:
@@
where $\psi$ corresponds to a message function equipped with learnable weights that compute a message between node $u$ and its neighbour $v$. $\bm{x}_{u}$ and $\bm{x}_{v}$ are the respective node embeddings. $\mathcal{N}(u)$ is the neighbourhood of node $u$, $\bigoplus$ refers to an aggregation function that defines how messages are passed Bronstein et al., (2021) and $\sigma$ is an activation function. There exist various different implementations of message passing layers, the scheme shown above, being the most general one Bronstein et al., (2021).
@@
A simple form of message passing is the spatial graph convolution. Here, messages are the features of the neighbouring nodes transformed using a learnable weight matrix. The aggregation corresponds to the summation operation. Morris et al., (2019) for example propose such an intuitive formulation:
@@
with $\bm{W}_{u}$ and $\bm{W}_{v}$ being learnable weights matrices, also referred to as filters. Typically, the weights are shared across all nodes and neighbours. This follows the concept of parameter sharing in CNNs.
@@
A method commonly used to improve the performance of a GNN model is to equip the graph convolution with an attention mechanism. Such attention layers are a special case of general message passing Bronstein et al., (2021). Hereby, attention coefficients are learned for each connected pair of nodes. They are computed from the features of the neighbouring nodes and the target node and determine the influence a neighbour has on the target node. Veličković et al., (2018) give a definition of such an attention convolution that would extend Eq. 2 to:
@@
with $\bm{a}$ and $\bm{W}_{a}$ being weight vector and matrix respectively. $\sigma$ again refers to an activation function, Veličković et al., (2018) for example use ReLu. The ||||| | corresponds to the concatenation operation, so the transformed features of both nodes are concatenated before the attention mechanism $\bm{a}$ is applied to produce the attention coefficients. The coefficients are normalized using softmax to make them comparable across the neighbours of the target node.
@@
Here, $\bm{U}$ is the matrix of eigenvectors of the normalized graph Laplacian $\bm{L}=\bm{I}-\bm{D}^{-\frac{1}{2}}\bm{A}\bm{D}^{-\frac{1}{2}}$ and is determined by the eigendecomposition $\bm{L}=\bm{U}\bm{\Lambda}\bm{U}^{T}$. Convolution in the spectral domain then is simply an element-wise multiplication, after which, the convolved signal is transformed back into the graph domain:
[…]
Where $\bm{U}^{T}\bm{g}$ is the filter in the spectral domain. Since the normalized graph Laplacian $\bm{L}$ is composed of the degree matrix $\bm{D}$ and adjacency matrix $\bm{A}$ ($\bm{I}$ being the unit matrix), intuitively its eigenvectors and eigenvalues indicate the major directions of information diffusion through the graph.
The spatial and spectral formulations of GNNs are equivalent, but due to the high computational cost of spectral GNNs resulting from the eigendecomposition at each step, spatial GNNs are more commonly used in many practical applications. In physical systems such as power grids, however, they appear relatively popular.
@@
Since GNNs are differentiable functions, they can be trained just like ordinary NNs, i.e., using gradient descent, backpropagation, batches or mini-batches. Commonly used loss functions include the negative log-likelihood of softmax functions for node or graph level classification. For link prediction, pairwise node embedding losses such as cross entropy or Bayesian personalised ranking loss are common.
@@
Due to their specific functionality, GNNs suffer from oversmoothing, oversquashing and scalability problems. Oversmoothing describes the problem that with an increasing number of layers, the features in each node become increasingly similar. This problem can for instance be addressed by regularisation or normalisation. Oversquashing refers to the distortion of information from distant nodes and is difficult to handle Giovanni et al., (2024). Regarding scalability, a GNN can either train on the entire graph and thus keep the full graph in memory. This may not be feasible for very large graphs. Or it can process mini-batches in the form of subgraphs, which leads to exponentially growing computational complexity with respect to the number of GNN layers Ding et al., (2022). While many ways to tackle this problem have been proposed, most of them sampling-based Wu et al., 2022a , scalability still remains a challenge.
@@ 3.2.2 Other Architectures used in the analyzed Approaches
The architecture proposed by Hamilton et al., 2017a is a special case of spatial graph convolution based on sampling. Instead of aggregating the entire neighbourhood of a node in each layer, a fixed number of neighbours of the target node are randomly sampled. The neighbours are then aggregated using a permutation invariant function such as mean, max or sum. They propose a loss function that allows unsupervised training of GraphSage. It consists of two terms, one enforcing that nodes that are close in the input graph have similar embeddings and the other pushing apart the embeddings of two nodes that are far apart in the graph.
@@
The idea of Graph Capsule Networks proposed by Verma and Zhang, (2018) is to capture more informative local and global features. This is done by constructing a capsule vector that contains enough discriminative features to allow proper reconstruction. These vectors are constructed using a capsule function that maps the input features of neighbouring nodes to higher order statistics. In a simple form, this could mean that the resulting vector contains the neighbourhood’s mean, the standard deviation and statistics of higher orders, depending on the given dimensionality of the capsule vector. The second key component of graph capsule networks is the aggregation function which is based on the covariance of the graph as it provides information such as norms or angles between node features.
@@ 3.3 Reinforcement Learning
RL is a key branch of machine learning thatfocuses on training agents to make sequential decisions in dynamic environments to maximize cumulative rewards. It primarily uses the framework of Markov Decision Processs to model decision-making problems. An MDP is defined by the tuple $M=(S,A,R,T,\gamma,H)$, capturing essential elements of the agent’s interaction with its environment. Figure 3 shows the schematic interaction between agent and environment in the RL framework.
@@ 3.3.1 Model-free RL
Value-based learning estimates the value of state-action pairs to select actions that maximize these values, using the value function $V^{\pi}(s)$ to represent the expected cumulative reward of following policy $\pi$ from state $s$.
$V^{\pi}(s)=\mathop{\mathbb{E}}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t})\mid\pi,s_{0}=s\right]$ (8)
Similarly, the action-value function $Q^{\pi}(s,a)$ represents the expected sum of future discounted rewards, beginning from state $s$, executing action $a$, and subsequently adhering to a given policy $\pi.$
@@
where $p(s,a,s^{\prime})$ models the state transition dynamics. In value-based approaches, finding the optimal policy involves identifying the optimal value function $Q^{*}(s,a)=\max_{\pi}Q^{\pi}(s,a)$. Explicit solutions to the Bellman equation are possible only when the dynamics function is known Sutton et al., (1999), so approximation methods are typically used. In the following, we present two such approximation methods.
  • Q-Learning is an off-policy learning method Watkins and Dayan, (1992) that aims to derive an optimal policy by directly updating values in a Q-table, a lookup table where each entry $Q(s,a)$ estimates the expected cumulative reward for taking action $a$ in state $s$. Q-Learning approximates the optimal action value function $Q^{*}$ through the following iterative updates.
@@
    Here, the agent explores the environment with a behavior policy, updating the Q-table based on observed rewards and state transitions.
  • Deep Q-Networks use neural networks to approximate action-value functions in high-dimensional input spaces, minimizing the mean-squared error between current and target Q-values Mnih et al., (2015). It introduces techniques like replay buffers and target networks to enhance training stability. Variants such as Double Deep Q-Network (DDQN) Hasselt et al., (2016), dueling DQN Wang et al., (2016), and Rainbow Hessel et al., (2017) further improve performance by addressing overestimation and efficiency issues.
@@
Policy-based learning directly estimates policies without intermediate value functions. It optimizes parameterized policies $\pi_{\theta}(a\mid s,\theta)$ that specifies the probability of action $a$ given state $s$ and parameters $\theta$ in order to maximize expected cumulative rewards. Unlike value-based approaches, policy-based methods update parameters using gradient-based optimization and are suitable in managing continuous action spaces and stochastic policies Sutton and Barto, (2018).
@@
Here, $\mu(s)$ is the distribution under $\pi$, $q_{\pi}(s,a)$ is the action-value function, and $\nabla\pi_{\theta}(a\mid s,\theta)$ is the gradient of $\pi$ with respect to $\theta$. This gradient is shown below and can be computed as the product of the policy probability and the gradient of the logarithm of the policy.
$\nabla_{\theta}\pi_{\theta}(a\mid s,\theta)=\pi(s,a)\nabla_{\theta}\log\pi_{\theta}(s,a)$ (13)
where $\nabla_{\theta}\log\pi_{\theta}(s,a)$ is the scoring function. Thus, the update of the policy parameters proceeds in the direction of the gradient of the objective function: $\Delta\theta=\alpha\nabla_{\theta}J(\theta)$, where $\alpha$ is the learning rate.
Actor-Critic methods
Actor-Critic algorithms combine the strengths of value-based and policy gradient-based learning. The actor learns policies to maximize rewards, while the critic evaluates these policies by estimating the value function. This framework addresses the limitations of both approaches and is fundamental to various RL algorithms, including Asynchronous Advantage Actor-Critic (A3C) Mnih et al., (2016), Deep Deterministic Policy Gradient (DDPG) Lillicrap et al., (2015), Proximal Policy Optimization (PPO) Schulman et al., (2017), and Soft Actor-Critic (SAC) Haarnoja et al., (2018).
For instance, A3C updates policy and value networks asynchronously across multiple threads, using an advantage function to guide actions and entropy regularization to enhance exploration. Similarly, DDPG, tailored for continuous action spaces, simultaneously learns a state-action value function (critic) and a policy (actor), employing experience replay and target networks for stable learning. In addition, PPO uses a clipped surrogate objective for smooth policy updates, balancing exploration and exploitation, which makes it a popular choice in RL research. Furthermore, SAC combines actor-critic methods with entropy regularization, training a policy network and two Q-value networks concurrently to encourage diverse action exploration and reduce overestimation.
While established algorithms such as PPO have been extensively used, newer innovations are often overlooked, particularly in power grid control. For instance, Bigger, Better, Faster (BBF) Schwarzer et al., (2023) is an advanced method that addresses the challenge of scaling networks effectively with limited samples. It employs a 15-layer ResNet architecture with widened layers, a high replay ratio Fedus et al., (2020) with periodic network resets, and adaptive strategies like dynamic update horizon and discount factor schedules. BBF discards NoisyNets Fortunato et al., (2017) in favor of weight decay for regularization, and outperforms other state-of-the-art agents in both computational efficiency and task performance, making significant strides in optimizing DRL for constrained data environments.

### v3

@@ 3.1.1 Transmission grids
Transmission grids operate at high voltages ranging from 110 kV to 765 kV. The exact voltage level varies from country to country. It is, therefore, not possible to make a clear distinction between voltage levels. In Germany, for example, the 110 kV network is considered a high-voltage distribution grid. The purpose of transmission grids is the same across all countries. It transports large amounts of electricity over long distances that vary from a few hundred kilometers (e.g., connecting offshore wind power in Northern Germany with the West-German industry) to a few thousand (e.g., China’s cross-country interconnections).
Transmission grid operation aims to achieve at least N-1 secure operation. This means that, if one asset in the grid fails, the grid is not overloaded. Therefore, transmission grids are typically built in meshed structures with redundancies installed, e.g., multiple transformers and busbars at substations. The structure of transmission grids results in a highly complex system where optimization problems have to be solved on a large scale.
One measure to prevent grid congestion, such as line overloading, is re-dispatch, which refers to changing generator injections. Since generation and consumption in a power system have to be balanced at all times, the change of one generator set point has to result in the change of another. This can result in renewable generators being shut down and fossil fuel generators being ramped up, which is undesirable in terms of both cost and CO2 neutrality. Therefore, other means of flexibility are researched such as topology control. Controlling the switching state in substations, the topology of the grid can be modified, helping to reduce or even eliminate the need for re-dispatch. Optimizing the topology is a challenge itself, as it results in a Mixed-Integer Non-Linear Problem. Here, deep learning solutions such as RL can help Marot et al., (2020).
@@ 3.1.2 Distribution grids
Distribution grids operate at lower voltages, typically below 110 kV, down to 120 V and 230 V. They deliver power from transmission grids to customers, such as industrial consumers and households.
Distribution grids cover smaller areas, like a village, and often don’t follow the N-1 security criterion due to the lower impact of asset failures. Radial or open-loop structures are common.
Voltage volatility is higher due to changing generation or consumption, especially with increased photovoltaic generation. This can cause voltage fluctuations. Traditional voltage control uses regulated transformers, shunt capacitors, and voltage regulators Srivastava et al., (2023). With digitalization, voltage control options are expanding to include inverter-based technologies like smart PV inverters Howlader et al., (2020), vehicle-to-grid systems Gonzalez Venegas et al., (2021), and stationary batteries Stecca et al., (2020).
Given the number of distribution grids, optimized control strategies need scalable and robust solutions.
@@ 3.1.3 Grid models
In research, synthetic grid models are widely used as a common benchmark to test new concepts and compare results. Examples in literature include Babaeinejadsarookolaee et al., (2021); Meinecke et al., 2020a ; Meinecke et al., 2020b ; Strunz et al., (2014). In the transmission and distribution grid use cases mentioned above, the IEEE test cases are the most commonly used benchmarks, available in several power system calculation tools like MATPOWER Zimmerman et al., (2011), pandapower Thurner et al., (2018), and DIgSILENT PowerFactory DIgSILENT GmbH, (2024).
However, these benchmark cases are simplified cases of power systems that do not fully capture the complexities and challenges of power systems. These grid models are mostly simple bus branch models with loads and generators connected to the buses. More sophisticated assets like inverter-based generator controls, transformer tap changers, shunt elements, etc., are not common for such grid models but are common in real power grid operation. Therefore, methods benchmarked on such synthetic grids have to be treated with caution as the application in practice requires further considerations.
@@ 3.2 Graph Neural Networks
Figure 2: Left: Visualization of the general message passing scheme in GNNs (modelled after Bronstein et al., 2021a ) - The target node (orange) receives messages $m_{u}i$ from its neighbours and aggregates them. The messages can be constructed from the information of both the target and neighboring node depending on the message passing scheme. Right: Illustration of a GNN (modeled after Wu et al., (2020)) - The graph is input to the GNN layers, which compute node embeddings based on the messages from neighboring nodes. As indicated in orange, this is done for each node in the graph. After all embeddings are computed, an activation function is applied. This is repeated for a given number of layers. In the end, the GNN outputs a graph with new node features from which prediction can be made.
GNNs are designed to extract information from graph-structured data by applying multiple layers of graph convolutions. They can be interpreted as a generalization of Convolutional Neural Networkss to non-euclidean structured data. The general idea is to combine information from local regions of the inputs in a learnable way and to grow these local regions from layer to layer. In this way, CNN layers learn increasingly abstract features from the input data. CNNs perform extremely well on grid-structured data, such as images. However, many real-world phenomena involve relationships or complex dependencies that cannot be represented as regular grid structures without losing information. For example, in an image, every node (pixel) has the same number of neighbors, but in power grids, not every component is connected to the same number of power lines.
Graphs consist of an unordered set of nodes and edges, where the edges define the neighborhood of a node. Graphs can, therefore, be used to model complex relationships, such as neighborhoods of arbitrary size or multiple types of edges or nodes. They can have attributes that describe properties of nodes and edges, such as node features or the strength of an edge. This makes graphs a perfect model for many real-world applications, such as power grids. In addition, feed-forward neural networks working on Euclidean data typically treat nodes as independent samples. This means they neglect the relationships between nodes or stack them unsystematically into a vector. GNNs, on the contrary, make use of the information about node connectivity. They can solve all common learning tasks, i.e., classification, regression, and clustering, for entire graphs and at node- or edge-level.
@@ 3.2.1 Message Passing
GNNs update the embeddings of the graph nodes by repeatedly aggregating information of their neighborhoods in a learnable way. This general scheme is called message passing, as each node updates its embedding based on the messages it receives from its neighbors. The representation $\bm{h}_{u}$ of a target node $u$ generated by a general message passing layer is computed as:
@@
where $\psi$ corresponds to a message function equipped with learnable weights that compute a message between node $u$ and its neighbor $v$. $\bm{x}_{u}$ and $\bm{x}_{v}$ are the respective node embeddings. $\mathcal{N}(u)$ is the neighborhood of node $u$, $\bigoplus$ refers to an aggregation function that defines how messages are passed Bronstein et al., 2021b and $\sigma$ is an activation function. There exist various different implementations of message passing layers, the scheme shown above, being the most general one Bronstein et al., 2021b .
@@
A simple form of message passing is the spatial graph convolution. Here, messages are the features of the neighboring nodes transformed using a learnable weight matrix. The aggregation corresponds to the summation operation. Morris et al., (2019) for example propose such an intuitive formulation:
@@
with $\bm{W}_{u}$ and $\bm{W}_{v}$ being learnable weights matrices, also referred to as filters. Typically, the weights are shared across all nodes and neighbors. This follows the concept of parameter sharing in CNNs.
@@
A method commonly used to improve the performance of a GNN model is to equip the graph convolution with an attention mechanism. Such layers are a special case of general message passing Bronstein et al., 2021b where attention coefficients are learned for each connected pair of nodes. They are computed from the features of the neighboring nodes and the target node and determine the influence a neighbor has on the target node. Veličković et al., (2018) give a definition of such an attention convolution that would extend Eq. 2 to:
@@
with $\bm{a}$ and $\bm{W}_{a}$ being weight vector and matrix respectively. $\sigma$ again refers to an activation function, Veličković et al., (2018) for example use ReLu. The ||||| | corresponds to the concatenation operation, so the transformed features of both nodes are concatenated before the attention mechanism $\bm{a}$ is applied. The coefficients are normalized using softmax to make them comparable across the neighbors of the target node.
@@
Where $\bm{U}$ is the matrix of eigenvectors of the normalized graph Laplacian $\bm{L}=\bm{I}-\bm{D}^{-\frac{1}{2}}\bm{A}\bm{D}^{-\frac{1}{2}}$ and is determined by the eigendecomposition $\bm{L}=\bm{U}\bm{\Lambda}\bm{U}^{T}$. $\bm{U}^{T}\bm{g}$ is the filter in the spectral domain. Since the normalized graph Laplacian $\bm{L}$ is composed of the degree matrix $\bm{D}$ and adjacency matrix $\bm{A}$, intuitively the eigenvectors and eigenvalues indicate the main directions of information diffusion through the graph. A first-order approximation of the spectral graph convolution has been proposed by Kipf and Welling, (2016):
$\bm{H}^{\prime}=\sigma(\bm{D}^{-\frac{1}{2}}A\bm{D}^{-\frac{1}{2}}HW)$ (6)
where $H$ corresponds to the node feature matrix and W is a learnable weight matrix.
While spatial and spectral formulations of GNNs are equivalent, spatial GNNs are more commonly used in practice due to the high computational cost of spectral GNNs from eigendecomposition.. However, they are more common in physical systems.
@@
Since GNNs are differentiable functions, they can be trained just like ordinary NNs, i.e., using gradient descent, backpropagation, batches, or mini-batches. Commonly used loss functions include the negative log-likelihood of softmax functions for the node or graph-level classification. For link prediction, pairwise node embedding losses such as cross-entropy or Bayesian personalized ranking loss are common.
@@
Due to their specific functionality, GNNs suffer from oversmoothing, oversquashing and scalability problems. Oversmoothing describes the problem that the node features become increasingly similar as the number of layers increases. This problem can be addressed by regularisation or normalization. Oversquashing refers to the distortion of information from distant nodes and is difficult to handle Giovanni et al., (2024). Regarding scalability, a GNN can either train on the entire graph and thus keep the full graph in memory. This may not be feasible for very large graphs. Or it can process mini-batches in the form of subgraphs, which leads to exponentially growing computational complexity with respect to the number of GNN layers Ding et al., (2022). While many ways to tackle this problem have been proposed, most of them sampling-based Wu et al., 2022a , scalability still remains a challenge.
@@ 3.2.2 Other Architectures used in the analyzed Approaches
The architecture proposed by Hamilton et al., 2017a is a special case of spatial graph convolution based on sampling. Instead of aggregating the entire neighborhood of a node in each layer, a fixed number of neighbors of the target node are randomly sampled. The neighbors are aggregated using a permutation invariant function such as mean or max. GraphSage is trained in an unsupervised manner using a special loss function. It consists of two terms, one enforcing that nodes that are close in the input graph have similar embeddings, and the other pushing apart the embeddings of two nodes that are far apart in the graph.
@@
The idea of Graph Capsule Networks proposed by Verma and Zhang, (2018) is to capture more informative local and global features. This is done using a capsule vector that contains enough discriminative features to allow proper reconstruction. These vectors are constructed using a capsule function that maps the node features to higher-order statistics depending on the given dimensionality of the capsule vector. In a simple form, this could mean that the resulting vector contains the mean and standard deviation of the node’s neighborhood. The second key component of graph capsule networks is the aggregation function, which is based on the covariance of a graph and provides information such as norms or angles between node features.
@@ 3.3 Reinforcement Learning
RL is a key branch of machine learning that focuses on training agents to make sequential decisions in dynamic environments to maximize cumulative rewards. It primarily uses the framework of Markov Decision Processs to model decision-making problems. An MDP is defined by the tuple $M=(S,A,R,T,\gamma,H)$, capturing essential elements of an agent’s interaction with its environment. Fig.  3 shows the schematic procedure of a RL framework.
@@ 3.3.1 Model-free RL
Value-based learning estimates the quality of state-action pairs to select optimal actions, i.e., actions with maximum value. For this purpose, the action-value function $Q^{\pi}(s,a)$ represents the expected sum of future discounted rewards, beginning from state $s$, executing action $a$, and subsequently adhering to a given policy $\pi.$
@@
where $p(s,a,s^{\prime})$ models the state transition dynamics. In value-based approaches, finding the optimal policy involves identifying the optimal value function $Q^{*}(s,a)=\max_{\pi}Q^{\pi}(s,a)$. Explicit solutions to the Bellman equation are possible only when the dynamics function is known Sutton et al., (1999). Therefore, approximation methods are typically used. Here, we present two such methods.
  • Q-Learning aims to derive an optimal policy by directly updating values in a Q-table, a lookup table where each entry $Q(s,a)$ estimates the expected cumulative reward for taking action $a$ in state $s$ Watkins and Dayan, (1992). Q-Learning approximates the optimal action-value function $Q^{*}$ through the following iterative updates.
@@
    Here, the agent explores the environment with a behavior policy, updating the Q-table based on the discrepancy between the actually observed and the previously expected reward.
  • Deep Q-Networks uses neural networks to approximate action value functions in high-dimensional input spaces, minimizing the error between current and target Q-values Mnih et al., (2015). It uses two networks, one to select actions and another to compute target Q-values. The target network is periodically updated with the weights of the primary network to stabilize the training. The agent stores its experience in a replay buffer from which it draws samples to train the neural network. Variants such as Double Deep Q-Network (DDQN) Hasselt et al., (2016), dueling DQN Wang et al., (2016), and Rainbow Hessel et al., (2017) further improve performance by addressing overestimation and efficiency issues.
@@
Policy-based learning directly estimates policies without intermediate value functions. It optimizes parameterized policies $\pi_{\theta}(a\mid s,\theta)$ that specify the probability of action $a$ given state $s$ and parameters $\theta$ to maximize expected cumulative rewards. A policy can be any mapping from state to action, for example, a neural network. Unlike value-based approaches, policy-based methods update parameters using gradient-based optimization and are suitable for continuous action spaces and stochastic policies Sutton and Barto, (2018).
@@
Here, $\mu(s)$ is the distribution under $\pi$, $q_{\pi}(s,a)$ is the action-value function, and $\nabla\pi_{\theta}(a\mid s,\theta)$ is the gradient of $\pi$ with respect to $\theta$. The update of the policy parameters proceeds in the direction of the gradient of the objective function to be maximized: $\Delta\theta=\alpha\nabla_{\theta}J(\theta)$, where $\alpha$ is the learning rate.
Actor-Critic methods.
These algorithms combine the strengths of value-based and policy gradient-based learning. The actor learns policies to maximize rewards, while the critic evaluates these policies by estimating the value function. This framework addresses the limitations of both approaches and is fundamental to various RL algorithms, including Asynchronous Advantage Actor-Critic (A3C) Mnih et al., (2016), Deep Deterministic Policy Gradient (DDPG) Lillicrap et al., (2015), Proximal Policy Optimization (PPO) Schulman et al., (2017), and Soft Actor-Critic (SAC) Haarnoja et al., (2018).
A3C updates policy and value networks asynchronously through multiple agents, using an advantage function to reinforce better-than-average actions. Further, it applies entropy regularization to enhance exploration, i.e., try new actions rather than exploiting knowledge already gained. Similarly, DDPG, tailored for continuous action spaces, simultaneously learns a state-action value function (critic) and a policy (actor), employing experience replay and target networks. PPO uses a clipped surrogate objective for smooth policy updates, balancing exploration and exploitation, making it a popular choice in RL research. Finally, SAC combines actor-critic methods with entropy regularization, training a policy network and two Q-value networks concurrently to encourage diverse action exploration and reduce overestimation.
While established algorithms such as PPO have been used extensively, recent innovations are often overlooked, particularly in power grid control. For instance, Bigger, Better, Faster (BBF) Schwarzer et al., (2023) is an advanced method that addresses scaling neural networks in a sample-efficient way. It employs a ResNet architecture with widened layers, a high replay ratio Fedus et al., (2020) with periodic network resets, and adaptive strategies like dynamic update horizon and discount factor schedules. BBF discards NoisyNets Fortunato et al., (2017) in favor of weight decay for regularization, and outperforms state-of-the-art agents in both computational efficiency and performance, enhancing DRL for constrained environments.